OpenStudy (anonymous):

Peanuts worth $4/kg were mixed with almonds worth $6/kg to produce 12 kg of nuts worth $62. Represent the unknowns above using the variables x and y.

OpenStudy (anonymous):

so let, peanuts = x and almonds = y

OpenStudy (anonymous):

the idea with these kind of questions is to form two equations and work them out simultaneously, so we know that a mix of peanuts and almonds = 12kg, as such we can see that 12kg = a certain amount of peanuts and a a certain amount of almonds or in other words 12kg = a certain amount of peanuts + a a certain amount of almonds 12kg = x +y

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

awesome got it Thanks

OpenStudy (anonymous):

u still need to figure out the second equation :)

OpenStudy (anonymous):

is the other equation 4x+6y=12(62)

OpenStudy (anonymous):

i did the first one and now ill guide u through with the second one :)

OpenStudy (anonymous):

4x + 6y = 62 is the second equation :) because $62 = $4 times the amount of peanuts (x) + $6 times the amount of almonds y =)

OpenStudy (anonymous):

I see. Thanks

OpenStudy (anonymous):

ur welcome :)